# Biharmonic nonlinear vector field equations in dimension 44

### v2

@@ 1 Introduction and main results
Let $\phi\in\mathcal{K}$ be such that $\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))>0$. Due to the uniform convergence $G(u_{j}+\phi)-G(u_{j})\to G(u+\phi)-G(u)$ on $supp\phi$, we have that
$\int_{\mathbb{R}^{4}}G(u_{j}+\phi)$ $=\int_{supp\phi}G(u_{j}+\phi)+\int_{\mathbb{R}^{4}\setminus supp\phi}G(u_{j})$
$\geq\int_{supp\phi}(G(u_{j}+\phi)-G(u_{j}))>0,$
for $j$ large enough. Thus,
@@
Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that as $t\to 0^{+}$, we have
@@
On the other hand, it is clear that
$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|,\text{ as }t\to 0^{+},$ (21)
holds for a constant $C_{G}>0$ depending only on $G$, since we have the bound $|G(v)|\leq C_{G}|v|$, provided that $|v|\leq\epsilon$.
Now, choose $\phi\in\mathcal{K}$ such that $\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0$. Then, for small enough $t>0$, it follows from (20) and (21) that $\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0$. By Lemma 3, we conclude that
@@
Let $\phi_{n}=u\rho_{n}\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
@@
Let also $A_{n}=\{x\in\mathbb{R}^{4}:n\leq|x|\leq 2n\}$. We assume by contradiction that $L(\phi)=0$, for all $\phi\in\mathcal{K}$ such that $\phi=0$ on $\{u=0\}$. It is clear that $\phi_{n}=0$ on the set $\{u=0\}$. Thus, we have $L(\phi_{n})=L(u\rho_{n})=0$. That is,
$\int_{\mathbb{R}^{4}}\sum_{i,j=1}^{4}\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\Big(\frac{\partial u}{\partial x_{i}}\frac{\partial\rho_{n}}{\partial x_{j}}+\frac{\partial u}{\partial x_{j}}\frac{\partial\rho_{n}}{\partial x_{i}}\Big)+\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big)\\ +\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}=0.$
@@
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\cdot\frac{\partial u}{\partial x_{i}}\Big)\frac{\partial\rho_{n}}{\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\Big|\Big(u\cdot\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\Big)\frac{\partial^{2}\rho_{n}}{\partial x_{i}\partial x_{j}}\Big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{\infty}(\mathbb{R}^{4})}\to 0\text{, as }n\to\infty,$
for a constant $k_{\rho}>0$ depending on $\rho$. This implies that $D^{2}u\equiv 0$, and since $u$ is bounded, we get $u\equiv 0$, which is a contradiction. ∎
Lemmas 3.5 and 3.6 from [8] are applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 and 7 below their proofs.
@@
We have that $\frac{G(u)}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}$ in $\mathcal{D}^{\prime}$.
@@
Let $M=\|u\|_{L^{\infty}(\mathbb{R}^{4})}$, and let $G_{k}$ be a smoothing sequence for $G$ so that $G_{k}\to G$ uniformly on $B_{M}=\{v\in\mathbb{R}^{m}:|v|\leq M\}$, and $g_{k}=\nabla G_{k}\to g$ pointwise on $B_{M}\setminus\{0\}$. Moreover, we can ensure that the maps $g_{k}$ are uniformly bounded on $B_{M}$. It is clear that $\frac{\partial G_{k}(u)}{\partial x_{i}}=g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}$. Thus, given $\phi\in C^{\infty}_{0}$ we have $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G_{k}(u)=-\int_{\mathbb{R}^{4}}\phi\big(g_{k}(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, and passing to the limit we get $\int_{\mathbb{R}^{4}}\frac{\partial\phi}{\partial x_{i}}G(u)=-\int_{\mathbb{R}^{4}}\phi\big(g(u)\cdot\frac{\partial u}{\partial x_{i}}\big)$, since $\frac{\partial u}{\partial x_{i}}=0$ on the set $\{u=0\}$. ∎
@@
Let $\lambda$ be as in Lemma 6. Then, $u$ is a solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $W^{4,q}_{loc}$, $\forall q\in[1,\infty)$, satisfying $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$ and $\int_{\mathbb{R}^{4}}G(u)=0$.
@@
The linear functional $M(\phi)=\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)$ restricted to $C^{\infty}_{0}$ satisfies $|M(\phi)|\leq C_{G}\|\phi\|_{L^{1}}$ (cf. (23)), thus by the Riesz Representation theorem, there exists $h\in L^{\infty}(\mathbb{R}^{4})$, $h:\mathbb{R}^{4}\to\mathbb{R}^{m}$, such that $M(\phi)=\int_{\mathbb{R}^{4}}h\cdot\phi$, $\forall\phi\in C^{\infty}_{0}$. Actually, (23) implies that $\int_{\mathbb{R}^{4}}h\cdot\phi=0$, for all $\phi\in C^{\infty}_{0}$ with support in the open set $\{u\neq 0\}$. Consequently, $h$ vanishes on $\{u\neq 0\}$, and denoting by $\chi$ the characteristic function, we have $\lambda\Delta^{2}u+g(u)=h\chi_{u=0}$ in $\mathcal{D}^{\prime}$. Our next claim is that $\lambda>0$. Otherwise, if $\lambda=0$, then $g(u)\equiv 0$. Moreover, in view of Lemma 7, the continuous function $x\mapsto G(u(x))$ is constant. Since $\lim_{|x|\to\infty}G(u(x))=0$, it follows that $G(u)\equiv 0$ and $u\equiv 0$, which is a contradiction. Hence, it holds $\Delta^{2}u=\frac{1}{\lambda}(g(u)-h\chi_{u=0})$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. Finally, since $u\in L^{\infty}$, while $g(u)-h\chi_{u=0}\in L^{\infty}$, we know by Lemma 9 that $u\in W^{4,q}_{loc}$, $\forall q\geq 1$. Therefore, $\Delta^{2}u=0$ a.e. on $\{u=0\}$, and $h=0$ a.e. on $\{u=0\}$. That is, we have $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. In addition, Pohozaev identity (cf. Lemma 10) implies that $\int_{\mathbb{R}^{4}}G(u)=0$, while in view of the weak convergence $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$, we have $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\leq T$. Thus, $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$, and $u$ solves the minimization problem (10). ∎
Setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e. $\bar{u}$ still solves the minimization problem (10)). Next, let $v\in\mathcal{C}$ be a solution of (2) in $\mathcal{D}^{\prime}$ such that $v\neq 0$. Then, by Pohozaev identity it holds $\int_{\mathbb{R}^{4}}G(v)=0$, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$. That is, $\bar{u}$ has all the properties stated in Theorem 1.1. We also point out that if $g$ is locally Hölder continuous, then by elliptic regularity we obtain that $\bar{u}\in C^{4}(\mathbb{R}^{4})$, since the function $x\mapsto g(\bar{u}(x))$ is locally Hölder continuous.
@@
As we mentioned in the Introduction, given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(6)), one can see that $\mathcal{C}\subset L^{2}(\mathbb{R}^{4})$. Let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$. We compute $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, since $\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}=1$. Consequently, we have the inequality $\frac{\mu^{2}}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\geq T$, which is equivalent to (12). Next, assume that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$, where $\bar{u}$ is a ground state solution to (2). Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the right hand side of (12) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the left hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$. This is an equality case in (12). Conversely, if the equality in (12) holds for $u\in\mathcal{C}^{\prime}$, i.e. $K(u)=Te^{2V(u)}$ (where $K$ is defined in (7)), then we have $K(\mu u)=T\Leftrightarrow\mu=e^{-V(u)}$. Moreover, for this value of $\mu$ there holds $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, that is, $\mu u$ solves the minimization problem (10). Therefore, in view of Remark 3, there exists a unique $r>0$ such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$. To this end, we use (11) (which also holds in the vector case since we have $|\nabla|u||\leq|\nabla u|$, $\forall u\in H^{1}(\mathbb{R}^{4};\mathbb{R}^{m})$), combined with the inequality (cf. [16, Lemma 6.1]):
$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4}),u\neq 0.$ (24)
In view of Lemma 11, we have $\mathcal{C}^{\prime}\subset H^{2}(\mathbb{R}^{4})$, as well as $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
@@
This proves that $\frac{1}{2T}<\frac{1}{(\pi e)^{2}}$.

### v3

@@ 1 Introduction and main results
Let $\phi\in\mathcal{K}$. We first notice that the sequence $\{(|u_{j}+\phi|-\epsilon/2)^{+}\}$ is uniformly bounded in $L^{4}(\mathbb{R}^{4})$. Indeed, this follows from the inequality
$\int_{|u_{j}+\phi|>\epsilon/2}|u_{j}+\phi|^{4}$ $\leq\int_{|u_{j}|>\epsilon/2}|u_{j}+\phi|^{4}+\int_{supp\,\phi\cap\{|u_{j}|\leq\epsilon/2\}}|u_{j}+\phi|^{4},$
combined with (18), and the inclusion $\mathcal{K}\subset L^{4}(\mathbb{R}^{4})$. Moreover, one can see that the sequence $\{(|u_{j}+\phi|-\epsilon/2)^{+}\}$ is also uniformly bounded in $W^{1,4}(\mathbb{R}^{4})$. Therefore, there exists $k_{1}>0$, such that the integrals $\int_{\mathbb{R}^{4}}A(2k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})$ are uniformly bounded. Next, we check that the sequence $\{G(u_{j}+\phi)\}$ is uniformly integrable on $\mathbb{R}^{4}$. To prove this, we use the bounds $|G(u)|\leq k_{2}+A(k_{1}(|u|-\epsilon/2)^{+})$, and $|A(|u|)|^{2}\leq k_{3}+A(2|u|)$, holding in $\mathbb{R}^{m}$ for some constants $k_{2},k_{3}>0$. Then, given $\omega\subset\mathbb{R}^{4}$, with $\mu(\omega)<\infty$, we obtain
$\int_{\omega}|G(u_{j}+\phi)|$ $\leq k_{2}\mu(\omega)+\int_{\omega}A(k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})$
$\leq k_{2}\mu(\omega)+(\mu(\omega))^{\frac{1}{2}}\Big(\int_{\omega}|A(k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})|^{2}\Big)^{\frac{1}{2}}$
$\leq k_{2}\mu(\omega)+(\mu(\omega))^{\frac{1}{2}}\Big(k_{3}\mu(\omega)+\int_{\mathbb{R}^{4}}A(2k_{1}(|u_{j}+\phi|-\epsilon/2)^{+})\Big)^{\frac{1}{2}},$
which implies the uniform integrability of $\{G(u_{j}+\phi)\}$ on $\mathbb{R}^{4}$. Consequently, since $G(u_{j}+\phi)\to G(u+\phi)$ a.e. in $\mathbb{R}^{4}$, we deduce that $G(u+\phi)$ is integrable on $\omega=supp\,\phi$, as well as on $\mathbb{R}^{4}$. In addition, we have $\lim_{j\to\infty}\int_{supp\,\phi}G(u_{j}+\phi)=\int_{supp\,\phi}G(u+\phi)$. The same argument also proves that the sequence $\{G(u_{j})\}$ is uniformly integrable on $\mathbb{R}^{4}$, and that $\lim_{j\to\infty}\int_{supp\,\phi}G(u_{j})=\int_{supp\,\phi}G(u)$. Therefore, we have $\lim_{j\to\infty}\int_{\mathbb{R}^{4}}(G(u_{j}+\phi)-G(u_{j}))=\int_{\mathbb{R}^{4}}(G(u+\phi)-G(u))$.
Finally, assuming (26), $u_{j}+\phi\in\mathcal{C}$ satisfies $\int_{\mathbb{R}^{4}}(G(u_{j}+\phi)-G(u_{j}))>0$, as well as $\int_{\mathbb{R}^{4}}G(u_{j}+\phi)>0$, for $j$ large enough. Thus, it follows that
@@
Note.
We shall see in the proof of Lemma 4 that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$, $\forall\phi\in\mathcal{K}$.
@@
Let $\phi\in\mathcal{K}$. By dominated convergence, one can see that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$, and as $t\to 0^{+}$, we have
@@
Indeed, it is clear that $\lim_{t\to 0}\frac{G(u(x)+t\phi(x))-G(u(x))}{t}=g(u(x))\cdot\phi(x)$, provided that $u(x)\neq 0$. On the other hand, since $(|u|-\epsilon/2)^{+}+|\phi|\in W^{1,4}(\mathbb{R}^{4})$ (cf. (19)), there exists $k_{1}>0$, such that $\int_{\mathbb{R}^{4}}A(2k_{1}[(|u|-\epsilon/2)^{+}+|\phi|])<\infty$. Next, in view of the bound $|g(u)|\leq k_{2}+A(k_{1}(|u|-\epsilon/2)^{+})$ and the monotonicity of $A$, we deduce that
$\forall s,t\in[0,1]:\ |g(u+st\phi)|$ $\leq k_{2}+A(k_{1}(|u+st\phi|-\epsilon/2)^{+})$
$\leq k_{2}+A(k_{1}[(|u|-\epsilon/2)^{+}+|\phi|]).$
Consequently, we obtain the inequality
$\Big|\frac{G(u(x)+t\phi(x))-G(u(x))}{t}\Big|=\Big|\int_{0}^{1}\big(g(u(x)+st\phi(x))\cdot\phi(x)\big)ds\Big|\\ \leq k_{2}|\phi(x)|+A(k_{1}[(|u(x)|-\epsilon/2)^{+}+|\phi(x)|])|\phi(x)|,$
whose right hand side is an integrable function (since $A(k_{1}[(|u(x)|-\epsilon/2)^{+}+|\phi(x)|])\in L^{2}(supp\,\phi)$ in view of the bound $|A(|u|)|^{2}\leq k_{3}+A(2|u|)$). This proves that $g(u)\cdot\phi\in L^{1}(\mathbb{R}^{4})$ and (30) holds.
Our next claim is that
$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|+o(t),\text{ as }t\to 0^{+}.$ (31)
To justify this estimate, we notice that given $\phi\in\mathcal{K}\subset W^{1,4}(\mathbb{R}^{4})$, there exists $k_{4}>0$ such that $A(k_{4}|\phi|)\in L^{1}(\mathbb{R}^{4})$. In view of the bound $|G(u)|\leq C_{G}|u|+k_{5}A(k_{4}|u|)$, holding in $\mathbb{R}^{m}$ for a constant $C_{G}>0$ depending only on $G$, and a constant $k_{5}>0$ depending also on $k_{4}$ (cf. (7)), it follows that
$\left|\int_{u=0}G(t\phi)\right|\leq C_{G}t\int_{u=0}|\phi|+k_{5}\int_{u=0}A(k_{4}t|\phi|).$
Finally, using the fact that $A(k_{4}t|\phi|)\leq t^{4}A(k_{4}|\phi|)$ holds for $0<t<1$, we deduce (31).
Now, choose $\phi\in\mathcal{K}$ such that
$\int_{\mathbb{R}^{4}}g(u)\cdot\phi-C_{G}\int_{u=0}|\phi|>0.$
Then, for small enough $t>0$, it follows from (30) and (31) that
$\int_{\mathbb{R}^{4}}\big(G(u+t\phi)-G(u)\big)>0.$
By Lemma 3, we conclude that
@@
Let $\phi_{n}=\rho_{n}u\in\mathcal{K}$, where $\rho_{n}=\rho(x/n)$, and $\rho\in C^{\infty}_{0}(\mathbb{R}^{4})$ is a function such that $0\leq\rho\leq 1$ and
@@
Let also $A_{n}=\{x\in\mathbb{R}^{4}:n\leq|x|\leq 2n\}$. We assume by contradiction that $L(\phi)=0$, for all $\phi\in\mathcal{K}$ such that $\phi=0$ on $\{u=0\}$. It is clear that $\phi_{n}=0$ on the set $\{u=0\}$. Thus, we have $L(\phi_{n})=L(\rho_{n}u)=0$. That is,
$\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\rho_{n}+\sum_{i,j=1}^{4}\big(\partial^{2}_{ij}u\cdot\big(\partial_{j}\rho_{n}\,\partial_{i}u+\partial_{i}\rho_{n}\,\partial_{j}u\big)+\big(u\cdot\partial^{2}_{ij}u\big)\partial^{2}_{ij}\rho_{n}\big)=0.$
@@
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\big|\big(\partial^{2}_{ij}u\cdot\partial_{i}u\big)\partial_{j}\rho_{n}\big|\leq k_{\rho}\|D^{2}u\|_{L^{2}(A_{n})}\|\nabla u\|_{L^{4}(A_{n})}\to 0\text{, as }n\to\infty,$
$\lim_{n\to\infty}\int_{\mathbb{R}^{4}}\big|\big(u\cdot\partial^{2}_{ij}u\big)\partial^{2}_{ij}\rho_{n}\big|$ $\leq\frac{k_{\rho}}{n}\|D^{2}u\|_{L^{2}(A_{n})}\|u\|_{L^{4}(A_{n})}$
$\leq\frac{k_{\rho}}{n}\|D^{2}u\|_{L^{2}(A_{n})}\big(\|(|u|-\frac{\epsilon}{2})^{+}\|_{L^{4}(A_{n})}+\frac{\epsilon}{2}\mu(A_{n})^{\frac{1}{4}}\big)$
[…]
where $k_{\rho}>0$ is a constant depending on $\rho$. This implies that $D^{2}u\equiv 0$ i.e. $u$ is affine, and since $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$, we get $u\equiv 0$, which is a contradiction. ∎
Lemmas 3.5 from [9] is applicable to our case without modifications. For convenience of the reader we reproduce in Lemmas 6 below its proof.
@@
We have that
$\frac{\partial[G(u)]}{\partial x_{i}}=g(u)\cdot\frac{\partial u}{\partial x_{i}}\text{ in }\mathcal{D}^{\prime}.$ (34)
Note.
It is clear that $G(u)\in L^{1}(\mathbb{R}^{4})$. On the other hand, we shall see in the proof of Lemma 7 that $g(u)\cdot\frac{\partial u}{\partial x_{i}}\in L^{1}_{loc}(\mathbb{R}^{4})$, thus (34) makes sense.
@@
As in [9, Lemma 3.6], we choose a smoothing sequence $\{G_{j}\}\subset C^{\infty}(\mathbb{R}^{m})$ for $G$, satisfying the following properties:
  • (i)
    $G_{j}\to G$, as $j\to\infty$, pointwise on $\mathbb{R}^{m}$,
  • (ii)
    $g_{j}=\nabla G_{j}\to g$, as $j\to\infty$, pointwise on $\mathbb{R}^{m}\setminus\{0\}$,
  • (iii)
    $\forall k_{2}>0$, $\exists k_{3}>0$ such that the inequalities $|g_{j}(u)|\leq k_{3}+e^{k_{2}|u|^{\frac{4}{3}}}$, and $|G_{j}(u)|\leq k_{3}+e^{k_{2}|u|^{\frac{4}{3}}}$, hold on $\mathbb{R}^{m}$ for all $j$.
Next, given an arbitrary open ball $B\subset\mathbb{R}^{4}$, we can find a sequence $\{v_{n}\}\subset C^{\infty}(\mathbb{R}^{4})$, such that $v_{n}$ converges to $u$ in $W^{1,4}(B)$, while $v_{n}$ and $\nabla v_{n}$ converge almost everywhere respectively to $u$ and $\nabla u$. In view of the Trudinger inequality [1, §8.27], there exists a constant $k_{1}>0$ such that the functions $e^{k_{1}|v_{n}|^{\frac{4}{3}}}$ are uniformly bounded in $L^{1}(B)$. Choosing $k_{2}=\frac{k_{1}}{2}$, it follows from (iii) above, that the functions $g_{j}(v_{n})$ and $G_{j}(v_{n})$ are uniformly bounded in $L^{2}(B)$. Consequently, given a test function $\phi\in C^{\infty}_{0}(B)$ and $i\in\{1,\ldots,4\}$, one can see that the sequences $\{G_{j}(v_{n})\partial_{i}\phi\}_{n}$ and $\{(g_{j}(v_{n})\cdot\partial_{i}v_{n})\phi\}_{n}$ are uniformly integrable over $B$. Since there holds
$\int_{B}G_{j}(v_{n})\partial_{i}\phi=-\int_{B}(g_{j}(v_{n})\cdot\partial_{i}v_{n})\phi,$
by letting $n\to\infty$, we deduce form the Vitali convergence theorem that
$\int_{B}G_{j}(u)\partial_{i}\phi=-\int_{B}(g_{j}(u)\cdot\partial_{i}u)\phi.$
Finally, by letting $j\to\infty$, and using the uniform integrability over $B$ of the sequences $\{G_{j}(u)\partial_{i}\phi\}_{j}$ and $\{(g_{j}(u)\cdot\partial_{i}u)\phi\}_{j}$, as well as the fact that $\partial_{i}u=0$ a.e. on the set $\{u=0\}$, we conclude again from the Vitali convergence theorem that
$\int_{B}G(u)\partial_{i}\phi=-\int_{B}(g(u)\cdot\partial_{i}u)\phi.$
@@
Let $\lambda$ be as in Lemma 6. Then, $\lambda>0$ and $u$ is a distributional solution of the equation $\Delta^{2}u=\frac{1}{\lambda}g(u)$, belonging to $W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\in[1,\infty)$.
@@
The linear functional $M(\phi)=\int_{\mathbb{R}^{4}}g(u)\cdot\phi-\lambda L(\phi)$ restricted to $C^{\infty}_{0}$ satisfies $|M(\phi)|\leq C_{G}\|\phi\|_{L^{1}}$ (cf. (33)), thus by the Riesz Representation theorem, there exists $h\in L^{\infty}(\mathbb{R}^{4})$, $h:\mathbb{R}^{4}\to\mathbb{R}^{m}$, such that $M(\phi)=\int_{\mathbb{R}^{4}}h\cdot\phi$, $\forall\phi\in C^{\infty}_{0}$. Furthermore, (33) implies that $|\int_{\mathbb{R}^{4}}h\cdot\phi|\leq C_{G}\int_{u=0}|\phi|$, for all $\phi\in C^{\infty}_{0}$, and thus for all $\phi\in L^{1}(\mathbb{R}^{4})$. Consequently, $h$ vanishes on $\{u\neq 0\}$, and denoting by $\chi$ the characteristic function, we have $\lambda\Delta^{2}u+g(u)=h\chi_{u=0}$ in $\mathcal{D}^{\prime}$. Our next claim is that $\lambda>0$. Indeed, if we assume that $\lambda=0$, then we obtain successively $g(u)\equiv 0$, and $\nabla(G(u))\equiv 0$ (by Lemma 7), from which it follows that $G(u)\equiv 0$, since $G(u)\in L^{1}(\mathbb{R}^{4})$. In particular we would have either $u(x)=0$ or $|u(x)|\geq\epsilon$, for a.e. $x\in\mathbb{R}^{4}$ (cf. (5)), and $\nabla(\min(|u|,\epsilon)\equiv 0$ in $\mathbb{R}^{4}$, which implies that the function $\min(|u|,\epsilon)$ is constant. This is excluded, since $u\in\mathcal{C}$ and $u\neq 0$. Hence, it holds $\Delta^{2}u=\frac{1}{\lambda}(g(u)-h\chi_{u=0})$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. Finally, we check that given $q>1$ and a bounded set $Q\subset\mathbb{R}^{4}$, we have $g(u)\in L^{q}(Q)$. To see this, we use the fact that $\int_{\mathbb{R}^{4}}A(k_{1}(|u|-\epsilon/2)^{+})<\infty$, for a constant $k_{1}>0$ (since $(|u|-\epsilon/2)^{+}\in W^{1,4}(\mathbb{R}^{4})$), together with the bounds $|g(u)|\leq k_{2}+A(\frac{k_{1}}{q}(|u|-\epsilon/2)^{+})$ and $|A(|u|)|^{q}\leq k_{3}+A(q|u|)$ holding in $\mathbb{R}^{m}$ for constants $k_{2},k_{3}>0$. Therefore, we have $u\in L^{q}_{loc}(\mathbb{R}^{4})$ as well as $g(u)-h\chi_{u=0}\in L^{q}_{loc}(\mathbb{R}^{4})$, $\forall q>1$, and in view of Lemma 9 we deduce that $u\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q>1$. This implies that $\Delta^{2}u=0$ a.e. on $\{u=0\}$, and $h=0$ a.e. on $\{u=0\}$. That is, we have $\Delta^{2}u=\frac{1}{\lambda}g(u)$ in $\mathcal{D}^{\prime}$, with $\lambda>0$. ∎
Finally, Pohozaev identity (cf. Lemma 10) implies that $\int_{\mathbb{R}^{4}}G(u)=0$, while in view of the weak convergence $D^{2}u_{j}\rightharpoonup D^{2}u$ in $L^{2}(\mathbb{R}^{4})$, we have $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\leq T$. Thus, $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}=T$, and $u$ solves the minimization problem (12). By setting $\bar{u}(x)=u(\lambda^{-\frac{1}{4}}x)$, we obtain a solution of (2), such that $\int_{\mathbb{R}^{4}}G(\bar{u})=0$, and $\frac{1}{2}\int_{\mathbb{R}^{4}}|D^{2}\bar{u}|^{2}=T$ (i.e. $\bar{u}$ still solves the minimization problem (12)). Next, let $v\in\mathcal{C}$ be a solution of (2) in $\mathcal{D}^{\prime}$ such that $v\neq 0$. Then, it holds $\int_{\mathbb{R}^{4}}G(v)=0$ in view of Pohozaev identity, and we have $S(v)=K(v)\geq T=K(\bar{u})=S(\bar{u})$. That is, $\bar{u}$ is a ground state solution. We also point out that the asymptotic convergence of $\bar{u}$ and $\nabla\bar{u}$ to 0 follows from Lemma 10.
@@
Given the potential $G(u)=|u|^{2}\ln|u|$ (which clearly satisfies assumptions (3)-(7)), we know by Remark 6 that $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$. Let $\mathcal{C}^{\prime}=\{u\in\mathcal{C}:\int_{\mathbb{R}^{4}}|u|^{2}=1\}$, let $u\in\mathcal{C}^{\prime}$, and set $\mu=e^{-V(u)}$, with $V(u)=\int_{\mathbb{R}^{4}}|u|^{2}\ln|u|$. We compute $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, since $\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}=1$. Consequently, we have the inequality $\frac{\mu^{2}}{2}\int_{\mathbb{R}^{4}}|D^{2}u|^{2}\geq T$, which is equivalent to (14). Next, assume that $u=\bar{u}/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$, where $\bar{u}$ is a ground state solution to (2). Then, setting $\mu=1/\|\bar{u}\|_{L^{2}(\mathbb{R}^{4})}$ one can see that the left hand side of (14) is equal to $\frac{1}{2}\ln(\mu^{2})$, while the right hand side is $V(\mu\bar{u})=\mu^{2}(\ln(\mu)\|\bar{u}\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(\bar{u}))=\ln(\mu)$, since $V(\bar{u})=0$. This is an equality case in (14). Conversely, if the equality in (14) holds for $u\in\mathcal{C}^{\prime}$, i.e. $K(u)=Te^{2V(u)}$ (where $K$ is defined in (9)), then we have $K(\mu u)=T\Leftrightarrow\mu=e^{-V(u)}$. Moreover, for this value of $\mu$ there holds $V(\mu u)=\mu^{2}(\ln(\mu)\|u\|^{2}_{L^{2}(\mathbb{R}^{4})}+V(u))=0$, that is, $\mu u$ solves the minimization problem (12). Therefore, in view of Remark 5, there exists a unique $r>0$ such that $\bar{u}(x)=\mu u(rx)$ is a ground state solution to (2).
To complete the proof of the theorem, it remains to show that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$. To this end, we use (13) (which also holds in the vector case since we have $|\nabla|u||\leq|\nabla u|$, $\forall u\in H^{1}(\mathbb{R}^{4};\mathbb{R}^{m})$), combined with the inequality (cf. [23, Lemma 6.1]):
$\int_{\mathbb{R}^{4}}|\nabla u|^{2}<\Big[\int_{\mathbb{R}^{4}}|\Delta u|^{2}\Big]^{\frac{1}{2}}\Big[\int_{\mathbb{R}^{4}}|u|^{2}\Big]^{\frac{1}{2}},\text{ for }u\in H^{2}(\mathbb{R}^{4};\mathbb{R}^{m}),u\neq 0.$ (35)
In view of Lemma 11, we have $\|D^{2}u\|_{L^{2}(\mathbb{R}^{4})}=\|\Delta u\|_{L^{2}(\mathbb{R}^{4})}$, thus we obtain
@@
This proves that $\frac{1}{2T}<\frac{1}{(2\pi e)^{2}}$.